OpenStudy (anonymous):

Simplify sqrt 48 . sqrt 75

jhonyy9 (jhonyy9):

4sqrt3 5sqrt3

jhonyy9 (jhonyy9):

there is multiplication ?

OpenStudy (anonymous):

yes, there is multiplication

OpenStudy (anonymous):

\[\sqrt{48} \times \sqrt{75} = \sqrt{48 \times 57} = \sqrt {3600} = 60\]